You are given a protein solution with a concentration of 0.15 mg/ml.

v. Suppose that we want to prepare 100 microliters of 10 micrograms/microliters solution. How much of H2O and protein stock do we need to add to obtain the target concentration and volume? (THE CONCENTRATION IS NOT GREATER IN THE QUESTION, YOU NEED TO CONVERT IT TO MICROGRAMS/MICROLITERS.... THEN IT SHOULD MAKE SENSE)

****IF YOU ARE GOING TO WRITE THIS ANSWER OUT WITH PEN AND PAPER PLEASE WRITE IN PRINT (NOT CURSIVE) AND LEGIBLY*****
Answer #1

Given that

Protein concentration of stock is 0.15mg/mL

This can be converted into microgram/microliters by multiplying with 10-3

Therefore the concentration of the stock protein is 0.15 x 10-3 microgram/microliter (N1)

Volume of stock protein (V1) to be taken to make final solution should be determined

Given that the concentration of the final solution is 10 microgram/microliter (N2)

Given that the volume of the final solution (V2) is 100 microliters

We have the formula that is N1V1 = N2V2

We have to determine V1 that is volume of stock protein solution to be taken

V1 = (N2V2)/N1 = (10 x 100) / 0.15 x 10-3 = 1/0.15 = 6.67 microliters

The volume of protein stock in the final 100 microliter solution is 6.67 microliters and the remaining is H2O

Therefore the volume of H2O = 100 - 6.67 = 93.33 microliters

Therefore we need 93.33 microliters of H2O (0.093 mL) and 6.67 microliters of stock protein (0.0067) is required to make our desired 100 microliter solution of 10 microgram/microliter solution.